# A DEIM-CUR factorization with iterative SVDs11footnotemark: 1

### v2

@@ 5 Error Analysis
In this section, we present a well-established theoretical error bound that pertains to a broad category of CUR factorizations. Consequently, this bound remains valid for the methods we propose. We refer the reader to [29, Section 4] for a comprehensive, constructive proof.
@@
[29, Thm. 4.1] Given $A\in\mathbb{R}^{m\times n}$ and a target rank $k$, let $V\in\mathbb{R}^{n\times k}$ and $U\in\mathbb{R}^{m\times k}$ be the leading $k$ right and left singular vectors of $A$. Suppose $C=AP$ and $R=S^{\top}\!A$ are of full rank, and $V^{\top}\!P$ and $S^{\top}U$ are nonsingular, then, with $M=C^{+}\!AR^{+}$, a rank-$k$ CUR decomposition constructed by the proposed techniques satisfies
@@ 6 Experiments
To assess the effectiveness of our algorithms, we test them on various data matrices from different application domains, such as economic modeling, categorizing text and retrieving information, and image compression. Our evaluation includes synthetic and real-world data matrices, both sparse and dense, with varying sizes ranging from small to large scale, which we summarize in Table 3. In the implementation, we use the in-built Matlab functions qr for the column pivoted QR and svds or svd for the SVD computation, where the latter is used for small-scale matrices. For Algorithms 6 and 7, we use our implementation of the Krylov–Schur method of [31] by incorporating the threshold parameter $\delta$ and upper limit $\ell$ on the number of singular vectors to be computed. Unless otherwise stated, in all the experiments we use as default the number of rounds $t=10$, the parameter $\delta=0.8$, and upper limit $\ell=k/10$.
@@ 6 Image compression Dense 150000 13072
From Fig. 2, we can see that in all cases our iterative subselection-based CUR algorithm has a lower approximation error than all the one-round deterministic index selection algorithms considered. We also observe that the approximation error of the QDEIM and the MaxVol techniques do not always decrease monotonically with increasing $k$ values. This phenomenon is local to the spectral norm; from results not presented here the errors in Frobenius norm are monotonically non-increasing. We choose the number of rounds for the CADP-CUR and CADP-CX algorithms to be $t=10$, and the parameter $\delta=0.8$ and upper limit $\ell=k/10$ for the DADP-CUR and DADP-CX algorithms. The results of all four proposed iterative subselection algorithms are comparable.
@@
As stated in Section 4, when dealing with large-scale matrices, performing a full (even reduced) SVD in each iteration of algorithms 3, 4, and 5 can often become excessively costly. To evaluate the efficiency of the proposed algorithms, which compute the full SVD, compared to their respective large-scale versions that utilize a Krylov–Schur routine to find a limited number of singular vectors (refer to Algorithm 7), we conduct experiments using a large sparse data set, i.e., the Reuters-21578. For our analysis, we approximate this matrix using a rank-50 approximation.

### v3

@@ 5 Error Analysis
In this section, we present a well-established theoretical error bound that pertains to a broad category of CUR factorizations. Consequently, this bound remains valid for the methods we propose. We refer the reader to [28, Section 4] for a comprehensive, constructive proof.
@@
[28, Thm. 4.1] Given $A\in\mathbb{R}^{m\times n}$ and a target rank $k$, let $V\in\mathbb{R}^{n\times k}$ and $U\in\mathbb{R}^{m\times k}$ be the leading $k$ right and left singular vectors of $A$. Suppose $C=AP$ and $R=S^{\top}\!A$ are of full rank, and $V^{\top}\!P$ and $S^{\top}U$ are nonsingular, then, with $M=C^{+}\!AR^{+}$, a rank-$k$ CUR decomposition constructed by the proposed techniques satisfies
@@ 6 Experiments
To assess the effectiveness of our algorithms, we test them on various data matrices from different application domains, such as economic modeling, categorizing text and retrieving information, and computational fluid dynamics. Our evaluation includes synthetic and real-world data matrices, both sparse and dense, with varying sizes ranging from small to large scale, which we summarize in Table 3. In the implementation, we use the in-built Matlab functions qr for the column pivoted QR and svds or svd for the SVD computation, where the latter is used for small-scale matrices. For Algorithms 6 and 7, we use our implementation of the Krylov–Schur method of [30] by incorporating the threshold parameter $\delta$ and upper limit $\ell$ on the number of singular vectors to be computed. Unless otherwise stated, in all the experiments we use as default the number of rounds $t=10$, the parameter $\delta=0.8$, and upper limit $\ell=k/10$.
@@ 5 Computational fluid dynamics Sparse 130412 30412
From Fig. 2, we can see that in all cases our iterative subselection-based CUR algorithms have lower approximation errors than all the one-round deterministic index selection algorithms considered. We also observe that the approximation error of the QDEIM and the MaxVol techniques do not always decrease monotonically with increasing $k$ values. This phenomenon is local to the spectral norm; from results not presented here the errors in Frobenius norm are monotonically non-increasing. We choose the number of rounds for the CADP-CUR and CADP-CX algorithms to be $t=10$, and the parameter $\delta=0.8$ and upper limit $\ell=k/10$ for the DADP-CUR and DADP-CX algorithms. The results of all four proposed iterative subselection algorithms are comparable.
@@
As stated in Section 4, when dealing with large-scale matrices, performing a full (even reduced) SVD in each iteration of algorithms 3, 4, and 5 can often become excessively costly. We evaluate the efficiency of the proposed algorithms: the small-scale versions, which compute the full SVD compared to their respective large-scale versions that use a Krylov–Schur routine to find a limited number of singular vectors (refer to Algorithm 7). This experiment uses a large sparse data set, i.e., the Reuters-21578. For our analysis, we approximate this matrix using a rank-50 approximation.